Qd=600-20px+0.02y-5pr
Qs=-300+10px

What are the equilibrium price and quantity?
1
Expert's answer
2019-03-18T11:21:26-0400

In equilibrium Qd = Qs, so:

600 - 20px + 0.02y - 5pr = -300 + 10px,

30px = 900 + 0.02y - 5pr,

px = 30 + 0.01/15y - 1/6pr,

Q = 0.1/15y - 5/3pr.